# Semi-Supervised In-Context Learning: A Baseline Study

### v1

@@ 3 Experimetnal Setup
Self-Consistency. Self-Consistency Wang et al. (2023) samples multiple responses using diverse decoding paths and selects the most consistent answer based on majority voting. The relative frequency of the majority answer naturally defines a confidence score for the generated annotations. Specifically, let $K_{i}$ be the number of times the response $y_{i}$ appears in the set of $K$ sampled responses. The Self-Consistency confidence score is then computed as:
$c_{\mathrm{SC}}=\max_{i}\frac{K_{i}}{K}.$ (6)
We discuss how the confidence scores are applied to each task in Appendix B.4.
Hyperparameters.
Unless stated otherwise, we filter all generated pseudo-demonstrations by the confidence at the 90th percentile. We discuss the hyperparameters of IterPSD in Appendix B.7.
Models.
We experiment with GPT-4o-mini and GPT-4o checkpointed on 2024-07-18 and 2024-11-20 on all of our experiments.
4 Empirical Analyses
4.1 Naive-SemiICL Consistently Beat Few-Shot Baselines
Figure 2 presents a comparison of the Few-Shot baseline and Naive-SemiICL (Naive). For Naive-SemiICL, We report the performance of the best performing confidence score [^4] A detailed breakdown of the performance across different confidence scores is provided in Appendix C.. Naive-SemiICL outperforms the baseline on all tasks, except on CLINC(D) under GPT-4o, where it ties with the baseline. To assess the impact of demonstration selection, we also examine the performance of Naive-SemiICL without the filtering step (Unfiltered). The unfiltered setting yields lower performance compared to Naive-SemiICL on all task except on CLINC(D) and FP under GPT-4o, and performs worse than the baseline in 20 out of 32 settings, which indicates that pseudo-demonstrations alone are insufficient to guarantee improvement over the baseline.
Overall, GPT-4o-mini benefits more from Naive-SemiICL, achieving a 9.94% performance gain over the baseline, while GPT-4o sees only a 6.05% improvement. This disparity may be attributed to GPT-4o’s stronger few-shot capabilities, as it outperforms GPT-4o-mini by an average of 4.93% in few-shot settings across all tasks.
@@ 4.2 IterPSD Improves Upon Naive-SemiICL
Table 1 demonstrates the superiority of IterPSD over Naive-SemiICL across five classification tasks using GPT-4o-mini. We evaluate both methods with Verbalized Confidence and Self-Consistency, with the best-performing hyperparameters detailed in Appendix B.7. IterPSD consistently outperforms Naive-SemiICL, particularly when paired with Self-Consistency, though the magnitude of improvement varies by dataset. Notably, IterPSD (Self-Consistency) achieves significant gains on BANKING, CLINC, CLINC(D) and FewEvent (over 3.0% performance gain), except for FP.
4.3 Scaling Law for Semi-Supervised ICL
Agarwal et al. (2024) identified a scaling law for In-Context Learning, where LLMs achieve peak performance with hundreds or thousands of ground-truth demonstrations. We observe a similar scaling law for Semi-Supervised ICL across classification and translation tasks. Figure 3 illustrates performance trends as a function of the number of sampled pseudo-demonstrations.
Across all configurations, Naive-SemiICL improves with larger demonstration sizes, though the point of peak performance varies. GPT-4o-mini scales effectively across most tasks, typically peaking between 500 and 1,000 examples, except for FP (100 examples) and Venetian (64 examples), where performance plateaus earlier. Similarly, GPT-4o reaches peak performance within the 500–1,000 range for classification tasks, with BANKING77 as an exception, achieving optimal performance with just 100 demonstrations. On translation tasks, GPT-4o exhibits a more stable scaling trend than GPT-4o-mini, with performance peaking later and declining more gradually.
We hypothesize that Naive-SemiICL’s decline in performance beyond a certain demonstration size stems from the accumulation of error in pseudo-demonstrations. To isolate the negative impact of long contexts on the LLMs, we examine the scaling behavior when all demonstrations are ground-truth data. Figure 5 presents results on six representative datasets, showing that both GPT-4o-mini and GPT-4o continue to improve beyond the peak of Naive-SemiICL. This suggests that the primary bottleneck lies in the quality of pseudo-demonstrations rather than the number of demonstrations.
4.4 Expanded Ground Truth Data
Is Semi-Supervised ICL still effective as the amount of ground-truth data increases, or does its effectiveness diminish with greater data availability? To answer this question, we evaluate Naive-SemiICL’s performance when $k_{g}\in\{32,64,100,500\}$ ground-truth demonstrations are available. We report the results using GPT-4o-mini with Verbalized Confidence on BANKING77, FewEvent and Bemba in Figure 4.
Across all three datasets, Naive-SemiICL outperforms most baselines, except for $k_{g}=100$ in FewEvent. The best performances are achieved with $k_{g}=500$, which performs on par with $k_{g}=64$ on FewEvent. We observe diminishing returns in performance gains as the number of annotated demonstrations increases. $k_{g}=64$ improves performance by 10.49% over the baseline, whereas $k_{g}=500$ yields only a 4.73% improvement on average across the three tasks. This indicates that increasing the number of ground-truth demonstrations improves performance, albeit with reduced gains. Additionally, we observe a scaling trend in BANKING77 and FewEvent similar to that reported in Section 4.3, but a generally downward trend on Bemba.
Figure 5: Many-shot scaling trends on classification and translation datasets. The blue solid line represents the performance of GPT-4o and GPT-4o-mini, while the gray dashed line indicates the number of demonstrations at which Naive-SemiICL achieves peak performance.
@@ 5 Related Work
Large Language Models (LLMs) exhibit remarkable zero-shot capabilities, allowing them to perform tasks without task-specific fine-tuning or prior examples Kojima et al. (2022). Their zero-shot predictions have proven to be effective sources of demonstration for in-context learning.
Auto-CoT Zhang et al. (2023) prompts the LLM with self-generated rationales on diversely sampled inputs. Rationales consisting of more than five reasoning steps are excluded from the demonstration to maintain the simplicity and accuracy of the demonstration. Such task-specific heuristic does not generalize to most recently published datasets such as LiveBench Math, as most of the generated rationale contains more than five steps. Our proposed Verbalized Confidence and Self-Consistency methods are task-agnostic and generally effective using the same cutoff threshold. Like Auto-CoT, Reinforced ICL Agarwal et al. (2024) generates rationales for reasoning problems and filters out those leading to incorrect answers. While this method requires the ground truth answer, our filtering method selects high-quality demonstration solely by the confidence score.
@@
Semi-supervised learning seeks to reduce reliance on labeled data by leveraging abundant unlabeled data to enhance model performance Lee et al. (2013); Sohn et al. (2020); Zou et al. (2025). Self-training McLachlan (1975); Xie et al. (2020) iteratively refines the model by using its own predictions on unlabeled data for training. Pseudo-labeling Lee et al. (2013); Sohn et al. (2020); Zou et al. (2023a, b) employs confidence-based filtering, retaining only high-confidence pseudo-labels to reduce error propagation and confirmation bias. JointMatch Zou and Caragea (2023) further alleviates error accumulation by using two independently initialized networks that teach each other through cross-labeling. Our work is the first to integrate confidence filtering and leverage both labeled and pseudo-labeled data in an in-context learning framework.
@@ 6 Conclusion
We introduced a semi-supervised In-Context Learning (ICL) framework that enhances self-generated pseudo-demonstrations through confidence-based data selection and iterative annotation. Our simple semi-supervised method, Naive-SemiICL, outperformed a strong 16-shot ICL baseline, achieving an average performance gain of 9.94% across 16 datasets. We further proposed IterPSD, an iterative refinement approach for pseudo-demonstrations, yielding up to 6.8% additional gains in classification tasks. Moreover, we identified a scaling law in semi-supervised ICL, showing that models achieve optimal performance with over 1,000 pseudo-demonstrations. Our analysis of Naive-SemiICL with increasing ground-truth data revealed diminishing returns—while additional ground-truth annotations improve performance, the relative contribution of pseudo-demonstrations decreases. This suggests that semi-supervised ICL is effective in low-resource settings, yet remains beneficial even with larger ground-truth datasets.
@@ B.1 Train-Test Split
B.2 Dataset Details
@@
Table 3: Best number of shots for the baseline on reasoning tasks.
B.3 Baseline Details
@@
B.4 Applying Confidence Scores
@@
B.5 Pseudo-Demonstration Sampling

### v2

@@ 4.2 Scaling Law for Semi-Supervised ICL
We hypothesize that Naive-SemiICL’s decline in performance beyond a certain demonstration size stems from the accumulation of errors in pseudo-demonstrations. To isolate the negative impact of long contexts on the LLMs, we examine the scaling behavior when all demonstrations are ground-truth data. Figure 5 shows that both GPT-4o-mini and GPT-4o continue to improve as the number of demonstrations increases, even beyond the optimal demonstration size for Naive-SemiICL in the 16-shot setting. This suggests that the performance degradation is not caused by long context length, but rather by the accumulated errors in pseudo-demonstrations. This finding motivates the design of IterPSD, which addresses error accumulation in pseudo-annotations through curriculum learning and iterative refinement.
Figure 4: Scaling trend of IterPSD on five benchmark tasks. Blue horizontal dashed line represents the best performing Naive-SemiICL on the same dataset.
@@ 4.3 IterPSD Improves Upon Naive-SemiICL
IterPSD outperforms Naive-SemiICL across five classification tasks, as shown in Table 3. We evaluate both methods using Verbalized Confidence and Self-Consistency. Notably, IterPSD achieves significant gains on BANKING, CLINC, CLINC(D), and FewEvent (over 3.0% performance gain), but not on FP. Similar to Naive-SemiICL, we observe a scaling law with respect to the number of pseudo-demonstrations used in IterPSD. Clear scaling trends are observed in four out of five tasks, as shown in Figure 4. On these tasks, IterPSD attains peak performance with 500 to 1,000 pseudo-demonstrations. The lack of scaling on FP may be attributed to the relative ease of the dataset, as Naive-SemiICL already achieved 98% accuracy on this task.
We also benchmark IterPSD on translation tasks, but the improvement over Naive-SemiICL is not consistent. We attribute this to the fact that each iteration of IterPSD needs to accumulate at least 100 demonstrations to avoid bias from sampling noise. However, Semi-Supervised ICL typically degrades after approximately 200 demonstrations, resulting in IterPSD terminating after 2 to 3 iterations.
@@ 5 Related Work
Large Language Models (LLMs) exhibit remarkable zero-shot capabilities, allowing them to perform tasks without task-specific fine-tuning or prior examples. Their zero-shot predictions have proven to be effective sources of demonstration for in-context learning Kojima et al. (2022); Zou et al. (2025a).
Auto-CoT Zhang et al. (2023) prompts the LLM with self-generated rationales on diversely sampled inputs. Rationales consisting of more than five reasoning steps are excluded from the demonstration to maintain the simplicity and accuracy of the demonstration. Such task-specific heuristic does not generalize to most recently published datasets such as LiveBench Math, as most of the generated rationales contain more than five steps. Li and Qiu (2023) builds on top of Auto-CoT with extra an extra step of semantic filtering. At each example during inference, the LLM is prompted to choose the demonstration for itself after retrieving the semantically relevant demonstrations through an embedding model. Like Auto-CoT, Reinforced ICL Agarwal et al. (2024) generates rationales for reasoning problems and filters out those leading to incorrect answers. While this method requires ground truths, our filtering method do so with self-generated confidence score.
@@
Semi-supervised learning seeks to reduce reliance on labeled data by leveraging abundant unlabeled data to enhance model performance Lee et al. (2013); Sohn et al. (2020); Zou et al. (2025b). Self-training McLachlan (1975); Xie et al. (2020) iteratively refines the model by using its own predictions on unlabeled data for training. Pseudo-labeling Lee et al. (2013); Sohn et al. (2020); Zou et al. (2023a, b) employs confidence-based filtering, retaining only high-confidence pseudo-labels to reduce error propagation and confirmation bias. JointMatch Zou and Caragea (2023) further alleviates error accumulation by using two independently initialized networks that teach each other through cross-labeling. Our work is the first to integrate confidence filtering and leverage both labeled and pseudo-labeled data in an in-context learning framework.
@@ 6 Conclusion
We introduced a semi-supervised ICL framework that enhances self-generated annotations through confidence-based data selection and iterative annotation. Our analysis of Naive-SemiICL with increasing amounts of ground-truth data reveals diminishing returns—while additional ground-truth annotations improve performance, the relative contribution of pseudo-demonstrations decreases. This suggests that semi-supervised ICL is particularly effective in low-resource settings, yet remains beneficial even when more ground-truth data is available. We further identify a scaling law in semi-supervised ICL, showing that models achieve optimal performance with over 1,000 pseudo-demonstrations. Our simple semi-supervised method, Naive-SemiICL, outperforms a strong 16-shot ICL baseline, achieving an average performance gain of 9.94% across 16 datasets. We also propose IterPSD, an iterative refinement approach for pseudo-demonstrations, which yields up to 6.8% additional gains on classification tasks.
@@ Appendix B Experimental Details
Figure 5: Many-shot scaling performance of GPT-4o-mini (top) and GPT-4o (bottom) across six selected datasets. The x-axis represents the number of shots (log scale), and the y-axis represents performance. The solid blue lines indicate many-shot in-context learning (ICL), while the dashed vertical lines mark the peak performance of Naive-SemiICL. Both models scale beyond the peak the performance of pseudo-demonstration approach.
@@ B.1 Train-Test Split
B.2 Computational Budget
We ran all of our experiments on an Apple M3 chip, where embedding-based search constitutes less than 1% of the computation time during IterPSD. The embeddings can be precomputed during data processing for each dataset, as it only needs to be computed once. It took about 400ms to retrieve each embeddings from the OpenAI API. The cost of generating the embeddings is \$0.13/million tokens. We ran all of our experiments on a \$1,000 budget.
B.3 Dataset Details
@@
Table 5: Best number of shots for the baseline on reasoning tasks. We use the same number of shots for Naive-SemiICL.
Model Bemba Fijian Faroese Venetian Tuvan Sardinian Banking FewEvent CLINC CLINCD FP
4o-mini 100 150 150 100 64 64 1000 2000 2000 2000 500
4o 100 150 100 200 100 100 1000 500 500 1000 100
Table 6: Optimal demonstration counts for Naive-SemiICL per dataset under 4o-mini and 4o models.
B.4 Baseline Details
@@
We experiment with different pseudo-demonstration sizes $k_{u}$ for Naive-SemiICL: $k_{u}\in\{32,64,100,150,200\}$ on classification tasks, and $k_{u}\in\{32,64,100,150,200\}$ on translation and reasoning tasks.
For MoT Li and Qiu (2023), we follow a recommended configuration of $5$ clusters. For retrieval, we employ the same text embeddings, text-embedding-3-large as Naive-SemiICL, and the same confidence threshold set at the 90th percentile. Since MoT needs to query the LLM $k$ times to select the most relevant examples, it is not suitable for classification and translation tasks that might utilize many examples, we only compare MoT to Naive-SemiICL on reasoning tasks.
B.5 Applying Confidence Scores
@@
B.6 Pseudo-Demonstration Sampling